# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 3. Quantum polymorphisms and reductions
Note that, if $A$ is finite (as we always assume), $A^{2}$ admits a family of $n=|A|^{2}$ generators. Indeed, we can simply let $g$ be the $i$-th projection $\mathbb{A}^{n}\to\mathbb{A}$, with $i\in[n]$ being the index of the pair $(c,d)$. However, $n$ could be smaller for certain $\mathbb{A}$. We point out that there are many cases of infinite structures $\mathbb{A}$ where $A^{2}$ is finitely generated under $\operatorname{Pol}(\mathbb{A})$; this is for example the case for the class of $\omega$-categorical structures, which has been widely studied in the context of classical constraint satisfaction problems.
@@
Let $\mathbb{A}$ be a structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$. Suppose also that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Then $\mathbb{A}^{n}$ is a commutativity gadget for $\mathbb{A}$.
@@
Every quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function. Using that $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ are generators for $A^{2}$, we know that, for each $i$, there exists $g_{i}\in\operatorname{Pol}(\mathbb{A})$ such that $g_{i}(\mathbf{a})=h_{i}(\mathbf{a})$ and $g_{i}(\mathbf{b})=h_{i}(\mathbf{b})$. Then $Q$ extends to a quantum homomorphism $Q^{\prime}:=g_{1}\oplus\cdots\oplus g_{d}$, thus proving ($(\mathbf{c_{1}})$). ∎
@@
Let $\mathbb{G}$ be an algebraic commutativity gadget with distinguished elements $u,v\in G$. Let $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$. We want to show that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, which is enough to conclude by Proposition 18.
Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that
@@
Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{q}\mathbb{A}^{n}$ given by the tensor product
@@
$U=Q\bullet W:\mathbb{G}\to_{q}\mathbb{A}.$
@@
The result is a direct consequence of Propositions 20 and 21 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 12). ∎
4. Warm-up: Quantum polymorphisms of cliques
@@
In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques. In particular, we prove the following.
@@
The proof goes by induction on $m\geq 3$.
For the case of $m=3$, it follows from [Ji, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 16 implies the result.
Take now $m\geq 4$. Let $Q\colon(\mathbb{K}_{m})^{n}\to_{q}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$. Pick $\mathbf{u},\mathbf{v}\in[m]^{n}$ and $i,j\in[m]$. Our goal is to show that $Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=0$.
@@
(3) $[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$
This is clearly the case if $Q_{\mathbf{w},k}=0$, so assume this is not the case. Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in\{1,\dots,n\}$, namely the subgraph induced by the neighborhood of $\mathbf{w}$. Note that $\mathbb{X}$ is isomorphic to $(\mathbb{K}_{m-1})^{n}$.
Define $R\colon\mathbb{X}\to_{q}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$. We check that this is a quantum homomorphism. Every $R_{\mathbf{z},\ell}$ is a projector since $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell}]=0$. For $\mathbf{z}\in X$, we have $\sum_{\ell}R_{\mathbf{z},\ell}=\sum_{\ell}Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}=Q_{\mathbf{w},k}(\sum_{\ell}Q_{\mathbf{z},\ell})=Q_{\mathbf{w},k}$, which is indeed the identity function on $Q_{\mathbf{w},k}(H)$ since $Q_{\mathbf{w},k}$ is a projector. Thus, $R$ is a quantum function. If $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$, then $\mathbf{z},\mathbf{z}^{\prime},\mathbf{w}$ form a triangle in $(\mathbb{K}_{m})^{n}$ and therefore the projectors $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell},Q_{\mathbf{z}^{\prime},\ell^{\prime}}$ pairwise commute for $\ell,\ell^{\prime}$ distinct from $k$. Thus, $($$\mathbf{QH_{2}}$)) holds. Finally, if $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$ and $\ell\neq k$, then $R_{\mathbf{z},\ell}R_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{w},k}Q_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{z}^{\prime},\ell}=0$. Thus, $($$\mathbf{QH_{1}}$)) holds.
@@
We can now already conclude by noting that
$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}],$
that all the terms with $k=\ell$ in the sum vanish by eq. 3, and that all the terms with $k\neq\ell$ in the sum vanish, since
@@
where the second equality is due to the fact that $Q_{\mathbf{w},k}$ and $Q_{\mathbf{w},\ell}$ commute with $Q_{\mathbf{u},i}$ and $Q_{\mathbf{v},j}$, and the third holds since $Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}=0=Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}$. ∎
Note that $\mathbb{K}_{2}$ has quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$. Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{q}\mathbb{K}_{2}$ by the following matrices:
@@
It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$. By Theorem 16, we obtain that $\mathbb{K}_{2}$ does not have a commutativity gadget.
@@
Let $m\geq 3$. Then $\mathbb{K}_{m}$ has a commutativity gadget. In fact, for every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
@@
A consequence of Proposition 23 is the undecidability of the entangled $m$-coloring problem for $m\geq 3$. The case of $m=3$ was proved in [culf_re-completeness_2025], and the case for $m>3$ was recently and independently proved in [Zeman, Proposition 5.2, Lemma 4.6]. We include a simple proof here using commutativity gadgets for the convenience of the reader.
@@
$\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
@@
We shall reduce from $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, whose undecidability was proved in [Ji, culf2024re]. Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, which we can assume does not contain isolated vertices. Let $\mathbb{Y}$ be the graph obtained by adjoining m3m-3 new vertices $y_{4},\dots,y_{m}$ to $\mathbb{X}$ and such that, whenever $x$ is a vertex in $\mathbb{X}$, then $\{x,y_{4},\dots,y_{m}\}$ forms a clique in $\mathbb{Y}$. Finally, adjoin a commutativity gadget between every $y_{i}$ and every other vertex in $\mathbb{Y}$. Such a commutativity gadget exists by Proposition 23. We let $\mathbb{Z}$ be the resulting structure.
If $\mathbb{X}$ has a quantum homomorphism $Q$ to $\mathbb{K}_{3}$, then $Q^{\prime}\colon\mathbb{Y}\to_{q}\mathbb{K}_{m}$ defined by $Q^{\prime}_{y_{i},i}=\operatorname{id}$ for all $i\in\{4,\dots,m\}$ and $Q_{x,j}=Q^{\prime}_{x,j}$ for all $x\in X$ is a quantum homomorphism. This $Q^{\prime}$ can be extended to a quantum homomorphism $\mathbb{Z}\to\mathbb{K}_{m}$ by ($(\mathbf{c_{1}})$).
Conversely, suppose that there exists a quantum homomorphism $Q\colon\mathbb{Z}\to_{q}\mathbb{K}_{m}$. By ($(\mathbf{c_{2}})$), it follows that for every $x\in X$, the restriction of $Q$ to $\{x,y_{2},\dots,y_{4}\}$ is non-contextual. Thus, this restriction can be written as $h_{1}\oplus\dots\oplus h_{d}$ for some classical homomorphisms $h_{1},\dots,h_{d}$. Up to composing the $h_{i}$ with an automorphism of $\mathbb{K}_{m}$, we can assume without loss of generality that $h_{i}(y_{j})=j$ for all $i\in\{1,\dots,d\}$ and $j\in\{4,\dots,m\}$. Thus, for each $i\in\{1,\dots,d\}$ and $x\in X$, we have that $h_{i}(x)\in\{1,2,3\}$. It follows that $\sum_{i=1}^{3}Q_{x,i}=\operatorname{id}$ and that $Q$ restricts to a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{K}_{3}$. ∎
5. Intermezzo: Contextuality and bifurcations
In this section, we present a general result on the emergence of a certain non-orthogonality structure in any contextual quantum homomorphism. While in the current work we shall only use this result to analyse quantum polymorphisms of odd cycles (see Section 6), we believe this result will be useful in a broader setting.
@@ 5. Intermezzo: Contextuality and bifurcations
Figure 1. A bifurcation of length $2$. Solid lines mean non-orthogonality.
The high-level idea is that, whenever some non-commutativity occurs in a quantum homomorphism $Q:\mathbb{X}\to_{q}\mathbb{Y}$, a bifurcation must appear in the product $X\times A$, in a sense that we formalise below. For simplicity, we restrict the analysis to binary structures (i.e., digraphs), although the same argument applies to quantum homomorphisms between arbitrary structures.

### v2

@@ 5.3 Complexity reductions via non-contextual quantum polymorphisms
Note that, if $A$ is finite (as is always the case in the current paper), $A^{2}$ admits a family of $n=|A|^{2}$ generators. Indeed, we can simply let $g$ be the $i$-th projection $\mathbb{A}^{n}\to\mathbb{A}$, with $i\in[n]$ being the index of the pair $(c,d)$. However, $n$ could be smaller for certain $\mathbb{A}$. This is for example the case for complete graphs (where $n=2$ is enough) or for undirected cycles (where $m$ is enough for an odd cycle of length $2m+1$). We point out that there are many cases of infinite structures $\mathbb{A}$ where $A^{2}$ is finitely generated under $\operatorname{Pol}(\mathbb{A})$; this is for example the case for the class of $\omega$-categorical structures, which has been widely studied in the context of classical constraint satisfaction problems.
@@
Let $\mathbb{A}$ be a (not necessarily finite) structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$. Suppose also that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Then $\mathbb{A}^{n}$ is a commutativity gadget for $\mathbb{A}$.
@@
Every quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to_{\operatorname{q}}A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function. Using that $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ are generators for $A^{2}$, we know that, for each $i$, there exists $g_{i}\in\operatorname{Pol}(\mathbb{A})$ such that $g_{i}(\mathbf{a})=h_{i}(\mathbf{a})$ and $g_{i}(\mathbf{b})=h_{i}(\mathbf{b})$. Then $Q$ extends to a quantum homomorphism $Q^{\prime}:=g_{1}\oplus\cdots\oplus g_{d}$, thus proving ($(\mathbf{c_{1}})$). ∎
@@
Let $\mathbb{G}$ be an algebraic commutativity gadget with distinguished elements $u,v\in G$. Let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$. We want to show that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, which is enough to conclude by Proposition 28.
Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that
@@
Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}^{n}$ given by the tensor product
@@
$U=Q\bullet W\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}.$
@@
The result is a direct consequence of Propositions 53 and 54 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 48). ∎
As a straightforward consequence of the above, we obtain the following result on reducibility between quantum CSPs.
Theorem 55.
Let $\mathbb{A},\mathbb{B}$ be relational structures such that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$ and $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$. Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
Proof.
The fact that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$ is equivalent to $\mathbb{A}$ admitting a commutativity gadget by Theorem 51, while the fact that $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$ is equivalent to $\mathbb{A}$ $p$-constructing $\mathbb{B}$ by Theorem 30. Using Theorem 50, we deduce that $\mathbb{A}$ q-constructs $\mathbb{B}$. Hence, there exists a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$ by Theorem 4, and the result follows from Theorem 3. ∎
We immediately derive from the above theorem the following useful undecidability criterion for quantum CSPs. As we shall see later, this generalises the known undecidability criteria for quantum CSPs.
Corollary 56.
Let $\mathbb{A}$ be a structure such that $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$ and all quantum polymorphisms of $\mathbb{A}$ are non-contextual. Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
Proof.
Since all quantum polymorphisms of $\mathbb{A}$ are non-contextual, we deduce that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Hence, by LABEL:{thm_comm_gadgets_plus_minion_homomorphisms_gives_reduction}, there is a logspace (in particular, computable) reduction from $\operatorname{CSP_{q}}(\operatorname{3SAT})$ to $\operatorname{CSP_{q}}(\mathbb{A})$. The result then follows from the undecidability of $\operatorname{CSP_{q}}(\operatorname{3SAT})$ [46, 60]. ∎
The result above can be rephrased in the following, convenient form.
See 7
Proof.
Since $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$, it follows from the algebraic CSP dichotomy theorem [21, 68] that $\mathbb{A}$ $p$-constructs $\operatorname{3SAT}$ and thus, by Theorem 30, $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$. To conclude, we apply Corollary 56. ∎
Note that Theorem 7 is in fact equivalent to Corollary 56 if $\mathsf{P}\neq\mathsf{NP}$. Indeed, under such hypothesis, $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$ if, and only if, $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$. On the other hand, if $\mathsf{P}=\mathsf{NP}$Theorem 7 becomes vacuous (unlike Corollary 56).
Example 57.
Recall Figure 1 in Section 2. The figure illustrates the undecidability of $\operatorname{CSP_{q}}(\mathbb{C}_{5})$ by reducing from quantum 1-in-3-SAT; i.e., from $\operatorname{CSP_{q}}(\mathbb{A})$ where $\mathbb{A}$ is the language encoding 1-in-3-SAT formulae. Formally, $\mathbb{A}$ is the Boolean structure with a single, ternary relation
$R^{\mathbb{A}}=\{(0,0,1),(0,1,0),(1,0,0)\}.$
The undecidability of quantum 1-in-3-SAT can be obtained by reducing from quantum 3SAT, using for example that $\operatorname{Pol}(\mathbb{A})$ does not contain the majority polymorphism, see [33] and Section 11.
5.4 Comparison with known undecidability criteria
The following undecidability criterion follows by combining [2, Lemma 10] with the undecidability of $\operatorname{CSP_{q}}(\mathbb{XOR})$ from [67], as pointed out in [63]. (For the formal definition of $\mathbb{XOR}$, see Section 11.)
Theorem 58 ([2, 67, 63]).
Let $\mathbb{A}$ be a Boolean relational structure such that $(i)$ its signature contains a binary symbol $R$ for which $R^{\mathbb{A}}=A^{2}$, and $(ii)$ $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$. Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
This result was generalised in [33] by replacing the requirement $(i)$ on the existence of a complete binary relation with a weaker structural condition called two-variable (non-)falsifiability, and relaxing the requirement on $\mathbb{A}$ being Boolean. Given any relation $S\subseteq A^{r}$ and two indices $i\neq j\in[r]$, we denote by $S_{i,j}$ the binary projection of $S$ onto the coordinates $i,j$; i.e.,
$S_{i,j}=\bigcup_{(a_{1},\dots,a_{r})\in S}\{(a_{i},a_{j})\}.$
Definition 59 ([33]).
Let $\mathbb{A}$ be a relational structure with signature $\sigma$. We say that $\mathbb{A}$ is two-variable falsifiable (TVF) if for each symbol $R\in\sigma$ and each pair $i\neq j\in[\operatorname{ar}(R)]$, it holds that $(R^{\mathbb{A}})_{i,j}\neq A^{2}$.
Theorem 60 ([33, Theorem 4.14]).
Let $\mathbb{A}$ be a relational structure such that $(i)$ $\mathbb{A}$ is not TVF, and $(ii)$ $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$. Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable. [^7] We point out that, just like for our Theorem 7, condition $(ii)$ in both Theorem 58 and Theorem 60 can be replaced by the algebraic condition $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$, which does not trivialise if $\mathsf{P}=\mathsf{NP}$.
Clearly, any structure satisfying the requirements of Theorem 58 also satisfies the requirements of Theorem 60, so the latter is indeed a generalisation of the former. We now prove that our undecidability criterion Theorem 7 extends the one in Theorem 60. To that end, we prove the following.
Proposition 61.
Let $\mathbb{A}$ be a relational structure. If $\mathbb{A}$ is not TVF, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Note that, in fact, Theorem 7 is strictly more general than Theorem 60. Indeed, there are multiple examples of structures $\mathbb{A}$ such that $(i)$ $\operatorname{CSP}(\mathbb{A})$ is NP-complete, $(ii)$ $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, but $(iii)$ $\mathbb{A}$ is TVF. For instance, as we shall see in later sections, the Siggers digraph, any odd cycle, and any clique give examples of such $\mathbb{A}$.
Proof of Proposition 61.
Let $\mathbb{A}$ be a $\sigma$-structure that is not TVF, and let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ be a quantum polymorphism of some arity $n\in\mathbb{N}$. Take two tuples $\mathbf{a},\mathbf{b}\in A^{n}$. We claim that the PVMs $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute. Since $\mathbb{A}$ is not TVF, there exists some symbol $R\in\sigma$ (of arity, say, $r$) and two indices $i\neq j\in[r]$ such that $R^{\mathbb{A}}$ has full binary projection onto the coordinates $i,j$; i.e., $(R^{\mathbb{A}})_{i,j}=A^{2}$. Therefore, we can assign to each $\ell\in[n]$ a tuple $\textbf{t}^{(\ell)}=(t_{1}^{(\ell)},\dots,t_{r}^{(\ell)})\in R^{\mathbb{A}}$ such that
$(t_{i}^{(\ell)},t_{j}^{(\ell)})=(a_{\ell},b_{\ell}).$ (8)
Let now $T$ be the $n\times r$ matrix whose rows are $\textbf{t}^{(1)},\textbf{t}^{(2)},\dots,\textbf{t}^{(n)}$. By definition of the categorical power $\mathbb{A}^{n}$, the $r$-tuple consisting of the list of columns of such matrix belongs to $R^{\mathbb{A}^{n}}$. This means that any pair of columns of $T$ having distinct indices is adjacent in $\operatorname{Gaif}(\mathbb{A}^{n})$. Note now that, because of (8), the tuples $\mathbf{a}$ and $\mathbf{b}$ occur in $T$ as the $i$-th and $j$-th columns, respectively. We deduce that $\mathbf{a}$ and $\mathbf{b}$ are adjacent in $\operatorname{Gaif}(\mathbb{A}^{n})$. Using $($$\mathbf{QH_{2}}$)), we conclude that $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ must commute, as claimed. ∎
We point out that Proposition 61 can alternatively be proved by observing that any relation having a full binary projection (and, thus, witnessing that $\mathbb{A}$ is not TVF) can be used to directly give a commutativity gadget for $\mathbb{A}$; then, the result follows from Theorem 51.
6 Quantum polymorphisms of cliques
@@
In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques. We let $\mathbb{K}_{m}$ denote the $m$-clique. We prove the following.
@@
The proof goes by induction on $m$.
For the case of $m=3$, it follows from [47, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 51 implies the result.
Take now $m\geq 4$. Let $Q\colon(\mathbb{K}_{m})^{n}\to_{\operatorname{q}}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$. Pick $\mathbf{u},\mathbf{v}\in[m]^{n}$ and $i,j\in[m]$. Our goal is to show that $Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=0$.
@@
$[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$ (9)
This is clearly the case if $Q_{\mathbf{w},k}=0$, so assume this is not the case. Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in[n]$, namely the subgraph induced by the neighborhood of $\mathbf{w}$. Note that $\mathbb{X}$ is isomorphic to $(\mathbb{K}_{m-1})^{n}$.
Define $R\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$. We check that this is a quantum homomorphism. Every $R_{\mathbf{z},\ell}$ is a projector since $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell}]=0$. For $\mathbf{z}\in X$, we have $\sum_{\ell}R_{\mathbf{z},\ell}=\sum_{\ell}Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}=Q_{\mathbf{w},k}(\sum_{\ell}Q_{\mathbf{z},\ell})=Q_{\mathbf{w},k}$, which is indeed the identity function on $Q_{\mathbf{w},k}(H)$ since $Q_{\mathbf{w},k}$ is a projector. Thus, $R$ is a quantum function. If $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$, then $\mathbf{z},\mathbf{z}^{\prime},\mathbf{w}$ form a triangle in $(\mathbb{K}_{m})^{n}$ and therefore the projectors $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell},Q_{\mathbf{z}^{\prime},\ell^{\prime}}$ pairwise commute for $\ell,\ell^{\prime}$ distinct from $k$. Thus, $($$\mathbf{QH_{2}}$)) holds. Finally, if $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$ and $\ell\neq k$, then $R_{\mathbf{z},\ell}R_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{w},k}Q_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{z}^{\prime},\ell}=0$. Thus, $($$\mathbf{QH_{1}}$)) holds.
@@
Observe now that
$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}].$
All terms with $k=\ell$ in the sum vanish by eq. 9. Moreover, all terms with $k\neq\ell$ in the sum vanish as well, since
@@
where the second equality is due to the fact that $Q_{\mathbf{w},k}$ and $Q_{\mathbf{w},\ell}$ commute with $Q_{\mathbf{u},i}$ and $Q_{\mathbf{v},j}$, and the third holds since $Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}=0=Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}$. This is enough to conclude. ∎
Note that $\mathbb{K}_{2}$ does have quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$. Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{\operatorname{q}}\mathbb{K}_{2}$ by the following matrices:
@@
It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$. By Theorem 51, we deduce that $\mathbb{K}_{2}$ does not have a commutativity gadget. In Section 11, we will see that this fact can be alternatively derived via a more general argument, by simply noting that $\mathbb{K}_{2}$ is preserved by the classical majority polymorphism and is not the full binary relation, see Theorem 11.
@@
Let $m\geq 3$. For every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
@@
An immediate consequence of Proposition 63 is the undecidability of the quantum $m$-coloring problem for $m\geq 3$.
@@
$\operatorname{CSP_{q}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
@@
The result follows by combining Theorem 62 and Corollary 56. ∎
We point out that the case of $m=3$ was proved in [33] via the reduction of [47], and the case of $m\geq 4$ was recently and independently proved in [34, Proposition 5.2, Lemma 4.6].
@@
Figure 3: A bifurcation of length $2$. Solid lines mean non-orthogonality.
7 Intermezzo: Contextuality and bifurcations
In this section, we present a general result on the emergence of a certain non-orthogonality pattern in any contextual quantum homomorphism. In later sections, we will use this result to analyse the quantum polymorphism minions of odd cycles and the Siggers digraph. We believe that the same ideas can be applied in a broader setting to prove non-contextuality of quantum polymorphisms (and, in particular, the existence of commutativity gadgets).
The high-level idea is that, whenever two PVMs do not commute in a quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$, a bifurcation must appear along some walk in the product $X\times A$, in a sense that we formalise below. Recall that the Gaifman graph $\operatorname{Gaif}(\mathbb{X})$ has vertex set $X$, and its edges are all pairs of vertices in $X$ simultaneously appearing in a tuple in some relation of $\mathbb{X}$.